This Question: 1 pt of 13 (4 This Quiz: 13 pts possi and standard deviation for the In a past election, the voter turnout was 79%, in a s numbers of voters in groups of 931. b. In the survey of 931 people, 693 said that they voted in the election. Is this result u kely to occur with a tumou of 79%? why or why not? c ased on these r results, does it appear that accurate voting results can be obtained by asking voters how ou they acted? a-?(Round to one decimal place as needed.) ?=?(Round to one decimal place as needed.) b. Is the result of 693 voting in the election usual or unusual? ou nn O B. This result is unusual because 693 is greater than the maximum usual O C. This result is unusual because 69 is within the range of usual values. OD. This resultis unusual because 693 is below the minimum usual value. c. Does it appear tuc that accurate voting results can be obtained by asking voters how they acted? did vote. 11:59pm earson Tutor
media%2F2ac%2F2ac860ef-f6ef-4d2b-b273-a4
Answer #1

a) \mu = n * p = 931 * 0.79 = 735.5

\sigma = sqrt(n * P * (1 - p))

    = sqrt(931 * 0.79 * 0.21)

    = 12.4

b) P(X = 693)

   = P(692.5 < X < 693.5)

   = P((692.5 - \mu )/\sigma < (X - \mu )/\sigma < (693.5 - \mu )/\sigma)

   = P((692.5 - 735.5)/12.4 < Z < (693.5 - 735.5)/12.4)

   = P(-3.47 < Z < -3.39)

   = P(Z < -3.39) - P(Z < -3.47)

   = 0.0003 - 0.0003

   = 0.00

As the probability value is less than 0.05, so it is unusul.

Option - d is correct.

c) Option - C